OpenStudy (anonymous):

Solve x-sqrt(1-x)=-5 algebraically.

OpenStudy (kinggeorge):

First up, move the square root over.\[x-\sqrt{1-x}=-5\]\[x+5=\sqrt{1-x}\]Now square both sides. Can you tell me what \((x+5)^2\) is?

OpenStudy (anonymous):

x^2+25? or is it x^2+10x+25?

OpenStudy (kinggeorge):

It's the second one. That means you have that \[x^2+10x+25=1-x\]Now you have the quadratic equation \[x^2+11x+24=0\]This factors easily. Can you factor this yourself as well?

OpenStudy (anonymous):

You are amazing!! Yes, I can factor that. It ends up being x+8 and x+3 yes? so x= -8, -3? but -8 will be extraneous, yes so the answer is -3?

OpenStudy (kinggeorge):

That looks perfect. Good job.